Question 1

In the economy of Zip, the marginal propensity to consume is 0.8. Investment is $60 billion, government expenditures on goods and services are $50 billion, and autonomous taxes are $60 billion. Zip has no exports and no imports.

(a)        The government increases its expenditures on goods and services to $60 billion. What is the change in equilibrium expenditure?

(b)        What is the value of the government expenditures multiplier?

(c)        The government continues to buy $60 billion worth of goods and services and increases autonomous taxes to $70 billion. What is the change in equilibrium expenditure?

(d)        What is the value of the autonomous tax multiplier?

(e)        The government simultaneously increases both its expenditures on goods and services and taxes by $10 billion. What is the change in equilibrium expenditure? Why does equilibrium expenditure increase?

Must accompany diagrams or visuals to aid the answer.
Answer #1

a.

MPC=0.8

MPS=1-MPC

=1-0.8

=0.2

Spending multiplier=1/MPS

=1/0.2

=5

If the government increases its expenditures on goods and services to $60 billion from $50 billion.

Change in the government spending=60-50

=$10 billion

Change in the equilibrium expenditure=$10 billion

b.

Government expenditure multiplier=1/MPS

=1/0.2

=5

c.

Tax multiplier=-MPC/MPS

=-0.8/0.2

=-4

Change in the tax=70-60

=10 billion

Change in the equilibrium income= change in the G* spending multiplier+ change in T* tax multiplier

=$10*5+ $10(-4)

=50-40

=$10 billion

d.

Tax multiplier=-MPC/MPS

=-0.8/0.2

=-4